# Rydberg superatoms: An artificial quantum system for quantum information processing and quantum optics

### v1

@@ III.1 Coherent laser excitation of Rydberg atom
Figure 5: Influence of Rabi frequency ($\Omega_{1}=\Omega_{2}$ for two-photon excitation schemes) on Rydberg excitation error in different schemes for an atomic temperature of (a) ${50}~{}{\mu}$K and (b) ${5.2}~{}{\mu}$K. The simulation is performed with intermediate detuning $\Delta_{e}$ = $2\pi\times 5687$ MHz in 780 nm - 480 nm scheme and $2\pi\times 7.6$ MHz in 420 nm - 1012 nm scheme.
@@
In addition, laser phase noise and intensity noise also contribute to the Rydberg excitation error. Depending on the laser phase stability and the Rabi frequency, the contribution of phase noise to the excitation error could be more than a few percent de Léséleuc et al. (2018b) without suppressing the servo bump. By using an optical cavity as a filter, as shown in Fig. 6(a), the servo bump, which is typically around MHz, is strongly suppressed by the cavity since its transmission bandwidth can be narrower than kHz. Levine et al. (2018) Moreover, Fig. 6(b) shows an alternative protocol, the feedforward technique, for suppressing the servo bump. A phase discrimination circuit based on the beat signal of the laser Li et al. (2022a) and the residual PDH error signal is used as the phase detector. Chao et al. (2023) The signal obtained from the phase detector is then feedforward to an electro-optical modulator (EOM) with optimized modulation depth so the phase noise of the laser can be reduced. An appropriate length of delay fiber is used to overcome the frequency deviations caused by the inherent delay time of the phase detector. It is reported that after suppressing the servo bump, the fidelity after $9\pi$ pulse  Li et al. (2022a) and the coherence time Levine et al. (2018) can be significantly increased. The intensity noise largely depends on the lasers themselves. For example, the 297 nm laser used in the single-photon scheme in ${}^{87}$b atoms can be obtained from the fourth harmonic generation seeded by a power-amplified 1188 nm laser. Using stimulated Raman scattering fiber amplifier technology, the amplified 1188 nm laser usually has a peak-to-peak intensity noise of a few percent. In addition, the cavities for frequency quadrupling will introduce additional intensity noise in the cavity locking. Considering the above noise sources, the resulting Rydberg excitation errors could be as large as a few percent.
III.2 Preparation of Rydberg superatom: Many-body Rabi oscillation
@@
A major difference between a Rydberg superatom and a single atom is their Rabi oscillations.Kumlin et al. (2023) For a single atom, if the laser field is resonant with the $|g\rangle\leftrightarrow|r\rangle$ transition, the system will undergo Rabi oscillations between the two states with Rabi frequency $\Omega$. Rabi et al. (1939) In 1954, Dicke predicted that the atom-light coupling strength could be increased by a factor of $\sqrt{N}$ in an ensemble.Dicke (1954) Such a collective effect could lead to a many-body Rabi oscillation with an enhanced Rabi frequency of $\sqrt{N}$$\Omega$.Cummings and Dorri (1983) In 2001, Lukin et al. suggested that if an atomic ensemble is confined within the blockade radius,Lukin et al. (2001) many-body Rabi oscillation could be achieved by performing laser excitation between the ground state |$\ket{G}$ and the Rydberg superatom state |$\ket{R}$. Coherent many-body Rabi oscillation is not only conclusive evidence for the successful preparation of the Rydberg superatom but also the cornerstone for many Rydberg-atom-related quantum applications, such as deterministic quantum photonic state preparation and manipulation, quantum networks, and quantum simulations.
@@
Figure 7: (a) Rydberg blockade between two single atoms. Reprinted with permission from Urban et al., Nature Physics 5(2), 110-114 (2009). Copyright 2009 Springer Nature.Urban et al. (2009) (b) Rabi oscillations enhanced by two atoms. Reprinted with permission from Gaëtan et al., Nature Physics 5(2), 115-118 (2009). Copyright 2009 Springer Nature.Gaëtan et al. (2009)
@@
Fig. 7(b) demonstrates the first two-atom enhanced ground-Rydberg Rabi oscillation.Gaëtan et al. (2009) To achieve a strong Rydberg blockade, two atoms are brought within a distance of 4 $\mu\rm m$. Careful selection of Rydberg levels results in the Förster resonance, leading to a Rydberg-Rydberg interaction strength $\sim$ 50 MHz. The Rabi oscillation of two atoms (blue squares) is accelerated compared to the case of a single atom (red circles), by a factor of 1.38(3), close to the expected value of $\sqrt{2}$ in the two-atom blockade model.
Figure 8: Many-body Rabi oscillations in atomic ensembles. (a) Coherent many-body Rabi oscillations of a mesoscopic atomic ensemble. Reprinted with permission from Dudin et al., Nature Physics 8(11), 790-794 (2012). Copyright 2012 Springer Nature.Dudin et al. (2012) (b) Rabi oscillation between the ground state and superatom state for various atom number distributions. Reprinted with permission from Ebert et al., Phys. Rev. Lett. 112(4), 043602 (2014). Copyright 2014 American Physical Society.Ebert et al. (2014) (c) Time evolution of the photon signal and the Rydberg population. Reprinted with permission from Paris-Mandoki et al., Phys. Rev. X 7(4), 041010 (2017). Copyright 2017 American Physical Society, licensed under a Creative Commons Attribution 4.0 International license.Paris-Mandoki et al. (2017)
@@
Several experiments conducted on atomic ensembles have also confirmed a similar many-body Rabi oscillation phenomenon. Fig. 8(b) shows the many-body Rabi oscillations where the number of atoms ranges from 1 to 15.Ebert et al. (2014) After the preparation of the Rydberg superatom |$\ket{R}$, it can be transferred to a ground state superatom |$\ket{G^{\prime}}$. The above protocol can be repeated to deterministically prepare the atomic Fock state |$\ket{G^{\prime}}^{n}$. Fig. 8(c) displays the many-body Rabi oscillation under strong dissipation,Paris-Mandoki et al. (2017) which facilitates the realization of a Rydberg-based single-photon absorber.Tresp et al. (2016)
Figure 9: Many-body Rabi oscillations in atom array experiments. (a) Collective Rabi oscillations in a quantum gas microscope. Reprinted with permission from Zeiher et al., Phys. Rev. X 5(3), 031015 (2015). Copyright 2015 American Physical Society, licensed under a Creative Commons Attribution 3.0 License.Zeiher et al. (2015) (b) Collective Rabi oscillation in two-dimensional atomic arrays. Reprinted with permission from Labuhn et al., Nature 534(7609), 667-670 (2016). Copyright 2012 Springer Nature.Labuhn et al. (2016) (c) Collective Rabi oscillation in one-dimensional atomic arrays. Reprinted with permission from Bernien et al., Nature 551(7682), 579-584 (2017). Copyright 2012 Springer Nature.Bernien et al. (2017)
In addition to atomic ensembles, many-body Rabi oscillations have also been observed in atom array settings. In 2015, Zeiher et al. demonstrated the preparation of a scalable Rydberg superatom in a quantum gas microscope [Fig. 9(a)].Zeiher et al. (2015) The $\sqrt{N}$ scaling in many-body Rabi oscillation was observed from one to a few hundred atoms. Related advances have also been made in reconfigurable tweezer arrays. Figs. 9(b) and 9(c) show the many-body Rabi oscillation with high contrast in one- and two-dimensional atomic arrays.Labuhn et al. (2016); Bernien et al. (2017)
@@ III.3 Quantum interface and coherence property
$\left|R\right\rangle=\frac{1}{\sqrt{N}}\sum_{j=1}^{N}e^{i\left(\mathbf{k}_{in}+\mathbf{k}_{c}\right)\cdot x_{j}}|g_{1},g_{2},...,r_{j},...,g_{N}\rangle\otimes|0\rangle.$ (17)
Here, $x_{j}$ is the position of the $j$-th atom, and $\mathbf{k}_{in},\mathbf{k}_{c}$ are the wave vectors of the input photon and the control laser, respectively. $|0\rangle$ is the vacuum state of the photon field. This Rydberg EIT light storage process efficiently converts a photon into a single Rydberg excitation collectively shared among the atoms inside the Rydberg blockade radius.
@@
$\left|E\right\rangle=\frac{1}{\sqrt{N}}\sum_{j=1}^{N}e^{i\left(\mathbf{k}_{in}+\mathbf{k}_{c}-\mathbf{k}_{r}\right)\cdot x_{j}}|g_{1},g_{2},...,e_{j},...,g_{N}\rangle\otimes|0\rangle,$ (18)
where $\mathbf{k}_{r}$ is the wave vector of the readout laser. Due to the finite lifetime in the intermediate state $\left|e\right\rangle$, the collective state $\left|E\right\rangle$ soon radiates into a new state,
$\frac{1}{\sqrt{N}}\sum_{j=1}^{N}e^{i\left(\mathbf{k}_{in}+\mathbf{k}_{c}-\mathbf{k}_{r}-\mathbf{k}_{out}\right)\cdot x_{j}}|g_{1},g_{2},...,g_{j},...,g_{N}\rangle\otimes|1\rangle_{\mathbf{k}_{out}},$ (19)
@@
Rydberg superatom also differs from a single atom in terms of its coherent properties. As a collective excitation, Rydberg superatom is the superposition of each microstate: $\sum_{j=1}^{N}e^{i\phi_{j}}|g_{1},g_{2},...,r_{j},...,g_{N}\rangle$. Therefore, the coherence of the Rydberg superatom largely depends on how well the relative phases $\phi_{j}$ between each microstates are maintained. Any effect that smears the relative phases will cause superatom decoherence.
@@
To suppress motional dephasing, an effective technique is to confine the atoms within spatial regions that are much smaller than the spin wave period. For ground states superatom, a one-dimensional optical lattice along the direction of the spin wave can be used to overcome the motional dephasing and extend the coherence time to minutes scale.Dudin, Li, and Kuzmich (2013) However, while red-detuned far-off-resonance optical traps are attractive for ground atoms, they are generally repulsive for Rydberg atoms. Therefore, traps that could simultaneously confine ground and Rydberg states are essential for extending motional decoherence times. To this end, optical traps with “magic” wavelengths are proposed Zhang, Robicheaux, and Saffman (2011); Saffman and Walker (2005) and demonstrated. Li, Dudin, and Kuzmich (2013) In ground-Rydberg state insensitive lattices, the Rydberg superatom storage times and many-body Rabi oscillation coherence times have been extended to tens of microseconds. Lampen et al. (2018); Mei et al. (2022)
@@
Due to their collective nature, the coherence properties between two Rydberg superatoms also differ from those of two single atoms. Multiple collective Rydberg excitations (multiple Rydberg superatoms) can be prepared for ensembles separated by a distance larger than the block radius. After the preparation of two Rydberg superatoms, the interaction between the pair of Rydberg atoms leads to the accumulation of phases over time,
$|RR\rangle\sim\sum_{j,j^{\prime}\neq j}^{N}e^{i\frac{V_{jj^{\prime}}}{\hbar}t}|\mathrm{~{}g}\rangle_{1}\ldots|r\rangle_{j}\ldots|r\rangle_{j^{\prime}}\ldots|\mathrm{g}\rangle_{N}.$ (21)
@@ IV APPLICATIONS IN QUANTUM INFORMATION PROCESSING
IV.1 Quantum logic gate
@@
Figure 11: (a) The displayed outcome presents both the results of Monte Carlo simulations and analytical approximations for a random gas in a finite volume, illustrating the probability distribution for frequency shifts of doubly excited Rydberg states. (b) Dipole blockade facilitates the generation of successive Fock states. Reprinted with permission from Lukin et al., Phys. Rev. Lett. 87(3), 037901 (2001). Copyright 2001 American Physical Society.Lukin et al. (2001)
@@
The strong and long-ranged interactions furnish a direct pathway for the realization of two-qubit gates within Rydberg superatoms. The pioneering gate protocol, as delineated in Ref. Lukin et al., 2001, entails a Rydberg superatom blockade gate. In an ensemble comprising $N$ identical atoms, each atom has the opportunity to be excited with equal probability, leading to the excitation of symmetric collective states exclusively starting from the ground state. However, the presence of resonant dipole-dipole interactions introduces a spectrum of doubly excited states, each state being separated by an energy gap of approximately $\hbar\bar{\kappa}$ in ensembles confined within a finite volume. Should the energy gap exceed the linewidth of the Rydberg state, resonant excitations from singly to doubly excited states experience significant suppression. Fig. 11 illustrates the probability distribution for frequency shifts of doubly excited Rydberg states by Monte Carlo simulations of $3\times 10^{4}$ positions within a rectangular box along with the corresponding analytical approximations for a random gas. The strength of the interaction $\kappa$ between any pair of atoms exceeds $\bar{\kappa}$, dominating most of the probability. This inhibition facilitates the production of collective spin states, enhancing the spectroscopic sensitivity beyond the standard quantum limit. It also aids in the preparation of a qubit that exhibits superpositions between the ground state and the single-excited storage state. This capability proves valuable for generating entanglement between two groups of atoms, drawing on principles similar to those described in Ref. Jaksch et al., 2000. This regime sets the groundwork for the emergence of superatoms in atomic gases, encompassing the Rydberg dressing, and opens avenues to scalable quantum logic gates within optical or magnetic trap arrays. Honer et al. (2010); Gaul et al. (2016); Jo et al. (2020) Subsequently, experimental demonstrations of superatom gates have been achieved, Reetz-Lamour et al. (2008) along with the proposal of a scheme combining electromagnetically induced transparency (EIT) and Rydberg blockade in an atomic ensemble. Weatherill et al. (2008)
@@
To streamline operational steps and address practical applications, a single-step multiatom entanglement scheme based on individual atoms within an ensemble has emerged. Müller et al. proposed a scheme to achieve a high-fidelity gate of many particles using Rydberg interactions and combining EIT, Müller et al. (2009) which can be applied to multiparticle interferometry measurements. Micheli et al. (2004); Seaman et al. (2007) In this work, a short $\pi$ pulse has no effect when the control atom is in the $|0\rangle$ state because it is decoupled from dynamics. The laser with Rabi frequency $\Omega_{c}$ between the intermediate state $|P\rangle$ and the Rydberg state $|R\rangle$ of the ensemble atoms is stronger than $\Omega_{p}$ between $|P\rangle$ and the ground states $|A\rangle$, $|B\rangle$, resulting in the EIT phenomenon when considering $\sqrt{2}\Omega_{p}/\Omega_{c}\ll 1$. The two ground states are decoupled, and the ensemble atoms always remain in the initial state $|A\rangle_{k}$. When the pulsed laser $\Omega_{p}$ is applied to the $k$th atom in the ensemble and undergoes adiabatic evolution along the dark state $|d\rangle_{k}$, causing decoupling of states $|A\rangle$ and $|B\rangle$ and achieving $|0\rangle|A^{N}\rangle$leftrightarrow $|0\rangle|A^{N}\rangle$. However, when the control atom initially stays at $|1\rangle$ and is excited to $|r\rangle$ by the laser with Rabi frequency $\Omega_{r}$, resulting in a shift of $|R\rangle$ because there is a blockade effect for the ensemble atoms so that $|P\rangle$ cannot couple with $|R\rangle$. Although $|A\rangle$, $|B\rangle$, and $|P\rangle$ remain non-resonantly coupled, $|1\rangle|B^{N}\rangle$leftrightarrow $|1\rangle|A^{N}\rangle$ can be achieved. Based on this model, the fidelity of the ${\rm C-NOT}^{N}$ logic gate can reach 99%. Similar structures can also be used for nonadiabatic geometric quantum computation to reduce evolution time and mitigate decoherence. Xu et al. (2012); Zhao et al. (2018); Maity and Purkait (2022); Guo et al. (2020) In 2018, nonadiabatic geometric quantum computation based on superatoms was proposed, Zhao et al. (2018) and can be extended to the 3-qubit case with two species of superatoms in a two-mode cavity. Maity and Purkait (2022) Guo et al. utilized dynamical invariant and ZZS methods to reverse design gate pulses, Guo et al. (2020) optimizing gate fidelity and robustness against control errors. The average fidelity of the two-qubit controlled-NOT gate in their scheme is 0.9971.

### v2

@@ III.1 Coherent laser excitation of Rydberg atoms
Figure 5: Influence of Rabi frequency ($\Omega_{1}=\Omega_{2}$ for two-photon excitation schemes) on Doppler-induced Rydberg excitation error in different schemes for an atomic temperature of (a) ${50}~{}{\mu}$K and (b) ${5.2}~{}{\mu}$K. The simulation is performed with intermediate detuning $\Delta_{e}$ = $2\pi\times 5687$ MHz in 780 nm-480 nm scheme and $2\pi\times 7.6$ MHz in 420 nm-1012 nm scheme.
@@
In addition, laser phase noise and intensity noise also contribute to the Rydberg excitation error. Depending on the laser phase stability and the Rabi frequency, the contribution of phase noise to the excitation error could be more than a few percent de Léséleuc et al. (2018b) without suppressing the servo bump. By using an optical cavity as a filter, as shown in Fig. 6(a), the servo bump, which is typically around MHz, is strongly suppressed by the cavity since its transmission bandwidth can be narrower than kHz. Levine et al. (2018) Moreover, Fig. 6(b) shows an alternative protocol, the feedforward technique, for suppressing the servo bump. A phase discrimination circuit based on the beat signal of the laser Li et al. (2022a) and the residual PDH error signal is used as the phase detector. Chao et al. (2023) The signal obtained from the phase detector is then fed forward to an electro-optical modulator (EOM) with optimized modulation depth so the phase noise of the laser can be reduced. An appropriate length of delay fiber is used to overcome the frequency deviations caused by the inherent delay time of the phase detector. It is reported that after suppressing the servo bump, the fidelity after $9\pi$ pulse  Li et al. (2022a) and the coherence time Levine et al. (2018) can be significantly increased. The intensity noise largely depends on the lasers themselves. For example, the 297 nm laser used in the single-photon scheme in 87Rb atoms can be obtained from the fourth harmonic generation seeded by a power-amplified 1188 nm laser.Li et al. (2019b) Using stimulated Raman scattering fiber amplifier technology, the amplified 1188 nm laser usually has a peak-to-peak intensity noise of a few percent. In addition, the cavities for frequency quadrupling will introduce additional intensity noise in the cavity locking. Considering the above noise sources, the resulting Rydberg excitation errors could be as large as a few percent.
III.2 Preparation of Rydberg superatoms: Many-body Rabi oscillation
@@
The concept of the Rydberg superatom can be connected to the original idea proposed by Dicke, where $N$ atoms are coupled to a single mode of the electromagnetic field at small interatomic distances. In the Dicke model, the collective behavior of closely spaced atoms leads to phenomenon such as superradiance, where the atoms act coherently as a single quantum system.Dicke (1954) This concept can be extended to the Rydberg superatom, where the interatomic distances can be much larger than the wavelength of the excitation light. Despite the larger distances, the atoms in a Rydberg superatom can still form collective excitation states, allowing them to behave as a single, coherent entity. Such a collective effect could lead to a many-body Rabi oscillation with an enhanced Rabi frequency of $\sqrt{N}$$\Omega$.Rabi et al. (1939); Cummings and Dorri (1983); Kumlin et al. (2023) In 2001, Lukin et al. suggested that if an atomic ensemble is confined within the blockade radius,Lukin et al. (2001) many-body Rabi oscillation could be achieved by performing laser excitation between the ground state |$\ket{G}$ and the Rydberg superatom state |$\ket{R}$ of Eq. (16). Coherent many-body Rabi oscillation is not only conclusive evidence for the successful preparation of the Rydberg superatom but also the cornerstone for many Rydberg-atom-related quantum applications, such as deterministic quantum photonic state preparation and manipulation, quantum networks, and quantum simulations.
@@
Figure 7: (a) Rydberg blockade between two single atoms. Reproduced with permission from Urban et al., Nature Physics 5(2), 110-114 (2009). Copyright 2009 Springer Nature.Urban et al. (2009) (b) Rabi oscillations enhanced by two atoms. Reproduced with permission from Gaëtan et al., Nature Physics 5(2), 115-118 (2009). Copyright 2009 Springer Nature.Gaëtan et al. (2009)
@@
Figure 7(b) demonstrates the first two-atom enhanced ground-Rydberg Rabi oscillation.Gaëtan et al. (2009) To achieve a strong Rydberg blockade, two atoms are brought within a distance of 4 $\mu\rm m$. Careful selection of Rydberg levels results in the Förster resonance, leading to a Rydberg-Rydberg interaction strength $\sim$ 50 MHz. The Rabi oscillation of two atoms (blue squares) is accelerated compared to the case of a single atom (red circles), by a factor of 1.38(3), close to the expected value of $\sqrt{2}$ in the two-atom blockade model.
Figure 8: Many-body Rabi oscillations in atomic ensembles. (a) Coherent many-body Rabi oscillations of a mesoscopic atomic ensemble. Reproduced with permission from Dudin et al., Nature Physics 8(11), 790-794 (2012). Copyright 2012 Springer Nature.Dudin et al. (2012) (b) Rabi oscillation between the ground state and superatom state for various atom number distributions. Reproduced with permission from Ebert et al., Phys. Rev. Lett. 112(4), 043602 (2014). Copyright 2014 American Physical Society.Ebert et al. (2014) (c) Time evolution of the photon signal and the Rydberg population. Reproduced with permission from Paris-Mandoki et al., Phys. Rev. X 7(4), 041010 (2017). Copyright 2017 American Physical Society, licensed under a Creative Commons Attribution 4.0 International license.Paris-Mandoki et al. (2017)
@@
Several experiments conducted on atomic ensembles have also confirmed a similar many-body Rabi oscillation phenomenon. Figure 8(b) shows the many-body Rabi oscillations where the number of atoms ranges from 1 to 15.Ebert et al. (2014) After the preparation of the Rydberg superatom |$\ket{R}$, it can be transferred to a ground state superatom |$\ket{G^{\prime}}$. The above protocol can be repeated to deterministically prepare the atomic Fock state |$\ket{G^{\prime}}^{n}$. Figure 8(c) displays the many-body Rabi oscillation under strong dissipation,Paris-Mandoki et al. (2017) which facilitates the realization of a Rydberg-based single-photon absorber.Tresp et al. (2016)
Figure 9: Many-body Rabi oscillations in atom array experiments. (a) Collective Rabi oscillations in a quantum gas microscope. Reproduced with permission from Zeiher et al., Phys. Rev. X 5(3), 031015 (2015). Copyright 2015 American Physical Society, licensed under a Creative Commons Attribution 3.0 License.Zeiher et al. (2015) (b) Collective Rabi oscillation in two-dimensional atomic arrays. Reproduced with permission from Labuhn et al., Nature 534(7609), 667-670 (2016). Copyright 2012 Springer Nature.Labuhn et al. (2016) (c) Collective Rabi oscillation in one-dimensional atomic arrays. Reproduced with permission from Bernien et al., Nature 551(7682), 579-584 (2017). Copyright 2012 Springer Nature.Bernien et al. (2017)
In addition to atomic ensembles, many-body Rabi oscillations have also been observed in atom array settings. In 2015, Zeiher et al. demonstrated the preparation of a scalable Rydberg superatom in a quantum gas microscope [Fig. 9(a)].Zeiher et al. (2015) The $\sqrt{N}$ scaling in many-body Rabi oscillation was observed from one to a few hundred atoms. Related advances have also been made in reconfigurable tweezer arrays. Figures 9(b) and 9(c) show the many-body Rabi oscillation with high contrast in one- and two-dimensional atomic arrays.Labuhn et al. (2016); Bernien et al. (2017)
@@ III.3 Quantum interface and coherence property
$\frac{1}{\sqrt{N}}\sum_{j=1}^{N}e^{i\left(\mathbf{k}_{in}+\mathbf{k}_{c}\right)\cdot\mathbf{x}_{j}}|g_{1},g_{2},...,r_{j},...,g_{N}\rangle\otimes|0\rangle.$ (17)
Here, $\mathbf{x}_{j}$ is the position of the $j$-th atom, and $\mathbf{k}_{in},\mathbf{k}_{c}$ are the wave vectors of the input photon and the control laser, respectively. $|0\rangle$ is the vacuum state of the photon field. This Rydberg EIT light storage process efficiently converts a photon into a single Rydberg excitation collectively shared among the atoms inside the Rydberg blockade radius.
@@
$\frac{1}{\sqrt{N}}\sum_{j=1}^{N}e^{i\left(\mathbf{k}_{in}+\mathbf{k}_{c}-\mathbf{k}_{r}\right)\cdot\mathbf{x}_{j}}|g_{1},g_{2},...,e_{j},...,g_{N}\rangle\otimes|0\rangle,$ (18)
where $\mathbf{k}_{r}$ is the wave vector of the readout laser. Due to the finite lifetime in the intermediate state $\left|e\right\rangle$, the collective state soon radiates into a new state,
$\frac{1}{\sqrt{N}}\sum_{j=1}^{N}e^{i\left(\mathbf{k}_{in}+\mathbf{k}_{c}-\mathbf{k}_{r}-\mathbf{k}_{out}\right)\cdot\mathbf{x}_{j}}|g_{1},g_{2},...,g_{j},...,g_{N}\rangle\otimes|1\rangle_{\mathbf{k}_{out}},$ (19)
@@
Rydberg superatom also differs from a single atom in terms of its coherent properties. As a collective excitation, Rydberg superatom is the superposition of each microstates: $\sum_{j=1}^{N}e^{i\phi_{j}}|g_{1},g_{2},...,r_{j},...,g_{N}\rangle$. Therefore, the coherence of the Rydberg superatom largely depends on how well the relative phases $\phi_{j}$ between each microstates are maintained. Any effect that smears the relative phases will cause superatom decoherence.
@@
To suppress motional dephasing, an effective technique is to confine the atoms within spatial regions that are much smaller than the spin wave period. For ground states superatom, a one-dimensional optical lattice along the direction of the spin wave can be used to overcome the motional dephasing and extend the coherence time to minutes scale.Dudin, Li, and Kuzmich (2013) However, while red-detuned far-off-resonance optical traps are attractive for ground atoms, they are generally repulsive for Rydberg atoms.Saffman and Walker (2005); Adams, Pritchard, and Shaffer (2020) Therefore, traps that could simultaneously confine ground and Rydberg states are essential for extending motional decoherence times. To this end, optical traps with “magic” wavelengths are proposed Zhang, Robicheaux, and Saffman (2011); Saffman and Walker (2005) and demonstrated. Li, Dudin, and Kuzmich (2013) In ground-Rydberg state insensitive lattices, the Rydberg superatom storage times and many-body Rabi oscillation coherence times have been extended to tens of microseconds. Lampen et al. (2018); Mei et al. (2022)
@@
Due to their collective nature, the coherence properties between two Rydberg superatoms also differ from those of two single atoms. Multiple collective Rydberg excitations (multiple Rydberg superatoms) can be prepared for ensembles separated by a distance larger than the blockade radius. After the preparation of two Rydberg superatoms, the interaction between the pair of Rydberg atoms leads to the accumulation of phases over time,
$|RR\rangle\sim\sum_{j,j^{\prime}\neq j}^{N}e^{i\frac{V_{jj^{\prime}}}{\hbar}t}|g_{1},...,r_{j},...,r_{j^{\prime}},...,g_{N}\rangle.$ (21)
@@ IV APPLICATIONS IN QUANTUM INFORMATION PROCESSING
IV.1 Quantum logic gates
@@
Figure 11: (a) The displayed outcome presents both the results of Monte Carlo simulations and analytical approximations for a random gas in a finite volume, illustrating the probability distribution for frequency shifts of doubly excited Rydberg states. (b) Dipole blockade facilitates the generation of successive Fock states. Reproduced with permission from Lukin et al., Phys. Rev. Lett. 87(3), 037901 (2001). Copyright 2001 American Physical Society.Lukin et al. (2001)
@@
The strong and long-range interactions furnish a direct pathway for the realization of two-qubit gates within Rydberg superatoms. The pioneering gate protocol in Ref. Lukin et al., 2001, entails a Rydberg superatom blockade gate. In an ensemble comprising $N$ identical atoms, each atom has the opportunity to be excited with equal probability, leading to the excitation of symmetric collective states exclusively starting from the ground state. However, the presence of resonant dipole-dipole interactions introduces a spectrum of doubly excited states, each state being separated by an energy gap of approximately $\hbar\bar{\kappa}=\wp_{rp^{\prime}}\wp_{rp^{\prime\prime}}/V$ in ensembles confined within a finite volume $V$, where $r$, $p^{\prime}$, and $p^{\prime\prime}$ denote long-lived Rydberg states, $\wp_{rp^{\prime}}$ and $\wp_{rp^{\prime\prime}}$ denote the dipole matrix elements for the corresponding transitions. Should the energy gap exceed the linewidth of the Rydberg state, resonant excitations from singly to doubly excited states experience significant suppression. Figure 11 illustrates the probability distribution for frequency shifts of doubly excited Rydberg states by Monte Carlo simulations of $3\times 10^{4}$ positions within a rectangular box along with the corresponding analytical approximations for a random gas. The strength of the interaction $\kappa$ between any pair of atoms exceeds $\bar{\kappa}$, dominating most of the probability. This inhibition facilitates the production of collective spin states, enhancing spectroscopic sensitivity beyond the standard quantum limit. It also aids in the preparation of a qubit that exhibits superpositions between the ground state and the single-excited storage state. This capability proves valuable for generating entanglement between two groups of atoms, drawing on principles similar to those described in Ref. Jaksch et al., 2000. This regime sets the groundwork for the emergence of superatoms in atomic gases, encompassing the Rydberg dressing, and opens avenues to scalable quantum logic gates within optical or magnetic trap arrays. Honer et al. (2010); Gaul et al. (2016); Jo et al. (2020) Subsequently, experimental demonstrations of superatom gates have been achieved, Reetz-Lamour et al. (2008) along with the proposal of a scheme combining EIT and Rydberg blockade in an atomic ensemble. Weatherill et al. (2008)
@@
To streamline operational steps and address practical applications, a single-step multiatom entanglement scheme based on individual atoms within an ensemble has emerged. Müller et al. proposed a scheme to achieve a high-fidelity gate of many particles using Rydberg interactions and combining EIT, Müller et al. (2009) which can be applied to multiparticle interferometry measurements. Micheli et al. (2004); Seaman et al. (2007). The basic idea is that a control atom conditionally affects the dynamics of spatially separated ensemble atoms through long-range Rydberg interactions. The control atom consists of two ground states $|0_{c}\rangle$ and $|1_{c}\rangle$, and a Rydberg state $|r_{c}\rangle$, where states $|1_{c}\rangle$ and $|r_{c}\rangle$ are resonantly coupled. Ground states $|A_{e}\rangle$ and $|B_{e}\rangle$ of atoms in the ensemble are off-resonantly coupled to intermediate state $|P_{e}\rangle$ (Rabi frequency $\Omega_{p}$), and to Rydberg state $|R_{e}\rangle$ (Rabi frequency $\Omega_{c}$). Under the EIT condition (two-photon resonance and $\Omega_{c}\gg\Omega_{p}$), the collective $\pi$-pulse evolution can be performed (e.g. $|0_{c}\rangle|A^{N}_{e}\rangle\to|0_{c}\rangle|A^{N}_{e}\rangle$) if the control atom is in state $|0_{c}\rangle$. However, when the control atom is initialized in state $|1_{c}\rangle$ and excited to $|r_{c}\rangle$, causing an energy shift of the $|R\rangle$ state. Using the resulted Raman transfer between $|A\rangle$, and $|B\rangle$, the collective operations $|1_{c}\rangle|A_{e}^{N}\rangle$leftrightarrow $|1_{c}\rangle|B_{e}^{N}\rangle$, and $|1_{c}\rangle|B_{e}^{N}\rangle$leftrightarrow $|1_{c}\rangle|A_{e}^{N}\rangle$ can be performed. Based on this model, the fidelity of the controlled-${\rm NOT}^{N}$ logic gate can reach 99%. Similar setting can also be used for nonadiabatic geometric quantum computation to reduce evolution time and mitigate decoherence. Xu et al. (2012); Zhao et al. (2018); Maity and Purkait (2022); Guo et al. (2020) In 2018, nonadiabatic geometric quantum computation based on superatoms was proposed, Zhao et al. (2018) and can be extended to the 3-qubit case with two species of superatoms in a two-mode cavity. Maity and Purkait (2022) Guo et al. utilized dynamical invariant and zero systematic-error sensitivity methods to reverse design gate pulses, Guo et al. (2020) optimizing gate fidelity and robustness against control errors. The average fidelity of the two-qubit controlled-NOT (CNOT) gate in their scheme is 0.9971.